OpenStudy (anonymous):

What is .2 equal to in a fraction?

OpenStudy (anonymous):

1/5 i think

OpenStudy (anonymous):

.2 = 2/10 = 1/5

OpenStudy (anonymous):

\[.2=\frac{1}{x}\] multiply by x \[.2x=1\] divide by .2 \[x=\frac{1}{.2}\] \[x=5\] \[.2=\frac{1}{5}\]

OpenStudy (anonymous):

thanks to all

OpenStudy (anonymous):

way is not 2/ 100?